What are three consecutive odd integers such that the first plus three times the second, minus twice the third is 44?
Found 2 solutions by rapaljer, mr_burn:
Answer by rapaljer(4671)

Let x= first integer
x+2 = second integer
x+4 = third integer
Now just do what the sentence says to do:
First + 3(Second) - 2(Third) = 44
x + 3(x+2) - 2(x+4) = 44
x + 3x + 6 - 2x - 8 = 44
2x -2 = 44
2x = 46
x=23 First
x+2 = 25 Second
x+4 = 27 Third
Check: Again, do what it says to do!
23 + 3(25) - 2(27)
23 + 75 - 54
98 - 54
44
It checks!

R^2 at SCC

Answer by mr_burn(1)

x= Three consecutive integers
x+3x-2x=44
2x=44
x=22
22+3(22)-2(22)=44
22+66-44=44
88-44=44
44=44
x=first odd integer
2x=second integer
3x=third integer
22=first integer
44=second integer
66=third integer